if p+sin 2 x+cos 4 x. then a)3\4 ≤ p ≤ 1(how ????)

if p+sin2x+cos4x.  then


a)3\4 ≤ p ≤ 1(how ????)


 

Grade:10

1 Answers

Badiuddin askIITians.ismu Expert
14 years ago

Dear ajinkya

p=sin2x+cos4x

   =1-cos2x + cos4x

   = 1-cos2xsin2x

   =1-1/4 sin22x

for maximum value of P , sin2x =0

P = 1

for minimum value of P  ,sin2x=1

P= 1-1/4

   =3/4

so    3/4 ≤ p ≤ 1

Please feel free to post as many doubts on our discussion forum as you can.
If you find any question Difficult to understand - post it here and we will get you the answer and

detailed
solution very  quickly.
 We are all IITians and here to help you in your IIT JEE  & AIEEE preparation.

 All the best.
 
Regards,
Askiitians Experts
Badiuddin